I ran 2.5 miles in 28 minutes, but I would like to know how many minutes it took to run 1.5 miles using the above information. Thank you.
Found 2 solutions by bulls400, jim_thompson5910:
Answer by bulls400(11)
Answer by jim_thompson5910(28598)

d = rt

2.5 = r(28)

2.5/28 = r

0.08928571 = r

r = 0.08928571


So the rate is roughly 0.08928571 miles per min. Now use this to find out how long it will take to run 1.5 miles


d = rt

1.5 = 0.08928571t

1.5/0.08928571 = t

16.8 = t

t = 16.8


So it will take 16.8 minutes.
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------